Suppose a soap-manufacturing production process is described by the following 

equation:

Y = a + b log K + с log L

Where, 

Y= Output (number of soaps produced)

K=Capital 

L=Labor

a, b and c are constants

Suppose 0<a<1, 0< b<1 

a. Find the Marginal Product of Labor (MPL) and Marginal Product of Capital (MPK)

in the production of soap

b. Is MPL diminishing, increasing or constant as L increases?

c. Is MPK diminishing, increasing or constant as K increases?


1
Expert's answer
2021-10-10T16:35:15-0400

a. The marginal product of labor (MPL) is calculated as follows:

"MP_L =\frac{ ∂Y}{∂L}\\=\frac{∂(a + b\space log\space K + c\space logL)}{∂L}\\=\frac{c}{L}"

The marginal product of labor (MPK) is calculated as follows:

"MP_K =\frac{ ∂Y}{∂K}\\=\frac{∂(a + b\space log\space K + c\space logL)}{∂L}\\=\frac{b}{K}"


b. The marginal product of labor is:

"MP_L =\frac{ c}{L}"

The labor is in denominator. If L increases then MPL decreases. Thus, it is diminishing.

c. The marginal product of capital is:

"MP_k =\frac{ b}{K}"

The capital is in denominator. If K increases then MPK decreases. Thus, it is diminishing.